OpenStudy (anonymous):

Hi, question. Does anyone know how to solve (5.4×10^-1)(7×10^0)

OpenStudy (whpalmer4):

there's no equation there to solve. Do you mean how to multiply it out?

OpenStudy (anonymous):

In scientific notation.

OpenStudy (anonymous):

I forgot to mention that detail.

OpenStudy (whpalmer4):

first, multiply the mantissa of one by the mantissa of the other. \[5.4*7 = \]

OpenStudy (anonymous):

Okay.

OpenStudy (whpalmer4):

what do you get?

OpenStudy (anonymous):

37.8

OpenStudy (whpalmer4):

good next, we multiply the two exponential parts: \[10^{-1}*10^0\]by adding the exponents:\[10^{-1}*10^{0} = 10^{-1+0} = 10^{-1}\] then we multiply that with the product we got for the mantissa part: \[37.8*10^{-1}\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Thank you

OpenStudy (whpalmer4):

Now, the mantissa is always supposed to have an absolute value between 1 and 10 if it isn't, then you need to normalize it

OpenStudy (whpalmer4):

which means divide or multiply by 10 as many times as needed, and adjust the exponent so that you aren't changing the value of the number.

OpenStudy (whpalmer4):

here we have 37.8, which is too big, so we divide by 10 to get 3.78, and we adjust the exponent up by 1. does that make sense?

OpenStudy (anonymous):

Not really. I'll probably have to go in before finals in the morning.

OpenStudy (whpalmer4):

we are essentially converting\[37.8*10^{-1}\]to\[\frac{1}{10}*37.8*10^{-1}*10\]to\[3.78*10^{-1+1} = 3.78*10^0\]

OpenStudy (anonymous):

Thank you

OpenStudy (whpalmer4):

37.8 * 0.1 is the same as 3.78 * 1, right? that's all we did in the normalization step

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Yup! I'd do the same haha

OpenStudy (whpalmer4):

such an answer deserves a medal :-)

OpenStudy (anonymous):

Okay. Thanks.